### v2

@@ 4 Data
In this section we discuss the characteristics and pre-processing of two distinguished and well-recognized datasets: the UNSW-NB15 [36] and CIC-IoT23 [37]. These two datasets were chosen due to their comprehensive nature and relevance in the field of cybersecurity, particularly in the context of network intrusion detection and IoT security. Both datasets are available from their official websites, and the pre-processing steps outlined below allow for full reproducibility.
4.1 UNSW-NB15
@@ 4.2 CIC-IoT23
5 Threat Detection Assumptions
@@
We assume access to active network traffic, specifically focusing on packet content from TCP and UDP transmission protocols. A significant portion of data transmissions occurs in plaintext [31, 32], which permits direct inspection of packet content. We also assume that analyzing the complete payload, even when it is fragmented across multiple packets, is sufficient for detection. With these assumptions, the proposed method offers several capabilities. By leveraging SSL during training, the model can learn robust representations of network bytes, enabling it to detect malware content early in active traffic. Furthermore, the method generalizes to unseen malware classes using few-shot learning, where limited labeled examples are sufficient to classify introduced attack types. The reliance on packet data also allows for efficient, real-time processing, making the approach scalable to high-throughput network traffic. Lastly, since the model assumes plaintext packet bytes, its effectiveness may diminish for encrypted traffic. To mitigate these risks, we evaluate the impact of encryption on detection performance by applying two symmetric encryption methods and repeating the method.
6 System Architecture and Model Training
In this section, we describe the system architecture, training, and evaluation of our model for network traffic classification. The overall architecture of the proposed method can be referred to in Fig. 3.
6.1 Transformer Encoder
@@
Embedding Layer: The embedding layer is the initial conversion phase between the input data and the transformer. The input, $\mathbf{X}=[x_{1},x_{2},\ldots,x_{N}]$, is transformed into a matrix representation, $\mathbf{E}\in\mathbb{R}^{N\times d_{\text{model}}}$, where $N$ is the number of decimal integer values in the sequence and $d_{\text{model}}$ is the dimension of the embedding vectors. The embedding layer converts these input sequences, encoded and padded during the pre-processing phase, into embedding vectors of fixed size 768. This conversion is performed by the learnable embedding matrix $\mathbf{W}_{E}$, which maps the fixed input tokens to their corresponding embedding vectors in the feature space. The output, $\mathbf{Z}$, is given by:
@@
The positional encoding vector is designed to have the same dimension, $d_{\text{model}}$, as the embedding layer, allowing the two to be summed together. The encoding for position $pos$ and dimension $i$ is calculated using sine and cosine functions:
@@
6.1.1 Input Linear Transformation
@@
6.1.2 Computation of Attention Scores
@@
6.1.3 Masked Prediction Layer
@@ 6.2 Model Training
19:  After convergence, freeze the transformer’s parameters $\theta$.
20:  For any input packet $p_{i}$, feed forward to obtain final-layer embeddings $\mathbf{E}_{i}$.
21:  Return the extract embeddings $\mathbf{E}$.
@@
18:     Compute embedding: $\mathbf{e}_{q}=f_{\phi}(q)$
@@
where $V$ is the number of unique packet elements, $y_{i,c}$ represents the ground truth for the $i$-th masked element, and $p_{i,c}$ is the predicted probability for each element $c$ at position $i$. $N$ is the total count of masked elements across the batch, and $N_{\text{masked}}$ is used to normalize the loss. This approach focuses the model on predicting masked elements using the context provided by the unmasked elements, leading to learn the representation of the data’s underlying structures.
The Adam optimizer is used with a learning rate of $52e-$ 5 and provides weight decay regularization, which prevents the weights from growing too large to help mitigate overfitting [39]. Overfitting occurs when a model fits too closely to the training dataset, leading to poor performance on the testing data. A scheduler for learning rate decay is also used to reduce the learning rate over the training period. The scheduler gradually increases the learning rate from zero to a specified learning rate during the warmup period, then linearly decreases the learning rate over the remaining epochs of training [40]. The experiments were conducted on NVIDIA GeForce RTX 2080 GPUs. This hardware enabled us to use significant computational power, facilitating quicker processing and learning.
6.3 Embedding Extraction
@@
6.4 Model Configuration
The selected hyperparameters of the model include the number of unique bytes (i.e. 258), hidden size, number of hidden layers, number of attention heads, the intermediate size, maximum length of the input payload, and the number of labels for this application, and are presented in Table 2.
The unique bytes parameter represents the total number of unique elements found in the input data. Here, the total is 258, accounting for each unique hexadecimal value and additional pad, unknown, and masking integers. The hidden size parameter is essential in determining the dimension of the hidden state. The number of hidden layers represents the number of stacked transformer blocks present in the model. Complementing these blocks is the number of attention heads, which control the model’s ability to attend to different segments of the input. This is pivotal for the model to comprehend and learn from the input data effectively. The intermediate size determines the dimension of the intermediate later in the network within each transformer layer. This value helps ensure that the network can process and transform the input data at each layer effectively. The maximum position embeddings denote the maximum length of the input sequence. For this study, we do not restrict the maximum packet length, which is set to 1500. It is important to note that transformer architectures are inherently capable of processing sequences that extend beyond a single packet’s MTU, meaning that if a payload exceeds 1460 bytes and is split across multiple packets, the transformer can process it as a continuous stream of bytes across packet boundaries. However, increasing the maximum input length significantly raises computational resource requirements, such as memory usage and processing time. These configurations remained consistent through our study of extracting the packet byte embeddings.
Table 2: Model Configurations
@@ 6.5 Downstream Task
In this architecture, the input embedding vector of dimension $d_{\text{model}}$ is first passed through a fully connected linear layer that projects it into a 512-dimensional space. This is followed by batch normalization, a ReLU activation to introduce non-linearity, and a dropout layer with a dropout rate of 0.5 to prevent overfitting. The network then reduces the 512-dimensional output to 128 dimensions through a second linear layer, again applying batch normalization and a ReLU activation to maintain stability and non-linearity in the learning process. Finally, a linear layer maps the 128-dimensional vector to $C$ output neurons, where $C$ corresponds to the number of classes. A softmax activation function is applied to the final output to generate a probability distribution over the classes.
@@
7 Few-Shot Learning Architecture
@@
In each episode, we randomly select $K$ classes from the dataset. For each class $k$, two distinct sets are sampled: a support set $S_{k}$ comprising $N_{S}$ examples, and a query set $Q_{k}$ comprising $N_{Q}$ examples. The algorithm computes the class prototypes $\mathbf{c}_{k}$ by taking the mean of the transformed feature vectors of the support set embeddings, $S_{k}$. Once the class prototypes are established, the algorithm evaluates each query sample $(x,y)$ in $Q_{k}$, calculating the Euclidean distance from $x$ to each class prototype $\mathbf{c}_{k}$, and then assigns the class label of the nearest prototype to $x$ The update to the model parameters is based on the loss $J$, which is formulated as the negative log-likelihood of the softmax probabilities over the distances between the query samples and the class prototypes. Specifically, for each query $(x,y)$ in $Q_{k}$, the loss $J$ is updated according to the equation:
$\scriptsize J\leftarrow J+\frac{1}{N_{C}N_{Q}}\left[d(f_{\phi}(x),c_{k})+\log\sum_{k^{\prime}}\exp(-d(f_{\phi}(x),c_{k^{\prime}}))\right].$ (10)
@@
Table 3: Results of Binary and Multiclass Classification of Malware Types on the Test Dataset.
Note: Bold font indicates the highest performance, while blue denotes the second-best result in each column.
@@
8 Results
In this section, we delve into the metrics to report performance, the state-of-the-art (SOA) benchmarks used for comparison, and the results of our proposed method.
8.1 Metrics
@@
8.2 Benchmarks
Table 4: Inference Time Per Sample (ms)
@@ 8.2 Benchmarks
Table 5: Results of Few-Shot Classification of Malware Types on the Query Set
@@
8.3 Binary and Multi-Class Classification
The data for the downstream classification task involving the SSL embeddings are divided into training, testing, and validation sets, with proportions of 70%, 20%, and 10%, respectively. The performance of our proposed method was evaluated over 25 epochs, demonstrating improvements in accuracy, precision, recall, and F1-Score when using packet bytes as input, shown in Table 3.
@@ 8.3 Binary and Multi-Class Classification
The inference times per sample, shown in Table 4, further emphasize the efficiency of the proposed method. The M.L. fully connected networks has a significantly lower inference time compared to other benchmark models. The M.L network achieves an inference time of 0.03 ms and 0.04 ms for UNSW-NB15 and CIC-IOT23, respectively. In contrast, models such as LSTM and CNNs exhibit considerably higher inference times, with the LSTM model requiring over 6 ms per sample on UNSW-NB15. The faster inference time of the M.L. network make it ideal for real-time applications.
@@
8.4 Few-Shot Classification
The primary motivation for these experiments is to assess the model’s ability to generalize to novel classes with limited supervision—a critical requirement for real-world scenarios where labeled data for all classes may not always be available. FSL allows us to test how well the SSL embeddings can adapt to previously unseen classes with very few labeled examples. To simulate this scenario, we train the SSL model exclusively on two of the three malware classes during pre-training, ensuring no prior exposure to the unseen class.
@@ 8.4 Few-Shot Classification
The results of our few-shot learning experiments for classifying both known and newly introduced malware types are presented in Table 5. The Trained Classes column indicates which classes were pre-trained with self-supervision, with the Unseen Class column serving as the novel, untrained class for that specific experiment. In the 1-shot setting on the UNSW-NB15 dataset, the model’s highest performance occurs when trained on Exploits and Generic malware types, with an overall query accuracy of 83.04% and 82.82% accuracy specifically for the unseen class. In the 5-shot setting, the best overall query accuracy of 89.48% is achieved when the model is trained on Fuzzers and Generic malware types, with Exploits as the unseen class. Meanwhile, the highest accuracy for an unseen class is observed for the Generic class, achieving 89.95%. For the CIC-IoT23 dataset, the model demonstrates a similar trend but with slightly lower accuracies, reflecting the increased diversity of the IoT traffic data. In the 1-shot setting, the best performance is achieved when training on Backdoor and Brute Force malware types, with Vulnerability as the unseen class, resulting in an overall query accuracy of 62.35% and 61.73% accuracy for the unseen class. Similarly, in the 5-shot setting, the model achieves its highest query accuracy of 75.01% when trained on Backdoor and Brute Force malware types, with Vulnerability as the unseen class.
@@
Lastly, we opt to compare the average few-shot results with the benchmarks to analyze how the accuracy compares, shown in Fig. 5. Despite using significantly fewer training samples, the FSL method performs comparably to SOA classifiers, surpassing methods like 2D-CNN, LSTM, and DEEPShield classifiers. Additionally, in our FSL framework, we train on only two of three classes, unlike the previously mentioned multiclass classification problem, which required all three classes to undergo self-supervised training. This in-turn saves training costs and allows for the ability to detect previously unseen malware classes with high accuracy, an issue that is overlooked in typical classification problems. Overall, our approach effectively leverages knowledge from known malware types to classify novel malware types that were unseen during training, even with a limited number of labeled samples. This capability makes it a promising solution for real-world scenarios where labeled data is scarce and new malware types are frequent.
8.5 Ablation Study
8.5.1 Masking Percentage
To investigate how varying proportion of masked bytes influences the quality of the learned embeddings, we conduct an ablations study examining multiple masking percentages ranging from 5% to 30%. Although 15% masking is generally used for SSL tasks, our objective is to learn the optimal ratio that preserves sufficient information for robust training of the model, but also prevents the model from becoming too reliant on unmasked data. Table 6 displays the results for different masking percentages on distinguishing between malicious and benign packets. On the UNSW-NB15 dataset, we can observe the highest accuracies occur between 5% and 15%, with accuracy slightly dropping off when masking more than 15% of packet bytes. On the more challenging CIC-IoT23 dataset, we observe that higher masking rates improve performance, with 25% byte masking resulting in over a 2% increase in classification over standard 15% masking.
Table 6: Masking Percentage Ablation with M.L. Classifier (Binary)

### v3

@@ IV Data
In this section, we discuss the characteristics and pre-processing of two distinguished and well-recognized datasets: UNSW-NB15 [37] and CIC-IoT23 [38]. These two datasets were chosen due to their comprehensive nature and relevance in the field of cybersecurity, particularly in the context of network intrusion detection and IoT security. Both datasets are available from their official websites, and the pre-processing steps outlined below allow for full reproducibility.
IV-A UNSW-NB15
@@ IV-B CIC-IoT23
V Threat Detection Assumptions
@@
We assume access to active network traffic, specifically focusing on packet content from TCP and UDP transmission protocols. A significant portion of data transmissions occurs in plaintext [31, 32], which permits direct inspection of packet content. We also assume that analyzing the complete payload, even when it is fragmented across multiple packets, is sufficient for detection. With these assumptions, the proposed method offers several capabilities. By leveraging SSL during training, the model can learn robust representations of network bytes without labels, enabling the detection of malware content early in active traffic. Furthermore, the proposed method generalizes to unseen malware classes using FSL, where limited labeled examples are sufficient to classify introduced attack types. The reliance on packet data also allows for efficient, real-time processing, making the approach scalable to high-throughput network traffic. Lastly, since the model assumes plaintext packet bytes, its effectiveness may diminish for encrypted traffic. To mitigate these risks, we evaluate the impact of encryption on detection performance by applying two symmetric encryption methods and repeating the method.
VI System Architecture and Model Training
In this section, we describe the system architecture, training, and evaluation of our model for network traffic classification. Our method operates in multiple stages: we first pre-train a byte-level transformer encoder with SSL through a MLM objective from scratch on unlabeled packet bytes, producing packet embeddings from the final hidden states of the transformer. Next, to validate the learned embeddings, we implement a lightweight downstream classifier on top of the frozen encoder for binary and multiclass malware detection and classification under the assumption that sufficient labeled data exists for conventional downstream classification. The overall architecture of the proposed method can be referred to in Fig. 3.
VI-A Transformer Encoder
@@
Embedding Layer: The embedding layer is the initial conversion phase between the input data and the transformer. The input, $\mathbf{X}=[x_{1},x_{2},\ldots,x_{N}]$, is transformed into the token-embedding matrix $\mathbf{Z}\in\mathbb{R}^{N\times d_{\text{model}}}$, where $N$ is the number of decimal integer values in the sequence and $d_{\text{model}}$ is the dimension of the embedding vectors. The embedding layer converts these input sequences, encoded and padded during the preprocessing phase, into embedding vectors of fixed size 768. This conversion is performed by the learnable embedding matrix $\mathbf{W}_{E}$, which maps the fixed input tokens to their corresponding embedding vectors in the feature space. The output, $\mathbf{Z}$, is given by:
@@
The positional encoding vector is designed to have the same dimension, $d_{\text{model}}$, as the embedding layer, allowing the two to be summed together. The encoding for position pospos and dimension $i$ is calculated using sine and cosine functions:
@@
VI-A1 Input Linear Transformation
@@
VI-A2 Computation of Attention Scores
@@
VI-A3 Masked Prediction Layer
@@ VI-B Model Training
19: After convergence freeze transformer’s parameters $\theta$ .
20: For any input packet $p_{i}$, feed forward to obtain last hidden state embeddings $\mathbf{H}_{i}$.
21: Compute the packet embedding $\mathbf{e}_{i}$ by mean pooling the rows of $\mathbf{H}_{i}$.
22:Return the set of packet embeddings $\{\mathbf{e}_{i}\}$.
@@
18:  Compute embedding: $\mathbf{e}_{q}=f_{\theta}(q)$
@@
where $V$ is the number of unique packet elements, $y_{i,c}$ represents the ground truth for the $i$-th masked element, and $p_{i,c}$ is the predicted probability for each element $c$ at position $i$. NN is the total count of masked elements across the batch, and $N_{\text{masked}}$ is used to normalize the loss. This approach focuses the model on predicting masked elements using the context provided by the unmasked elements, leading to learn the representation of the data’s underlying structures.
The Adam optimizer is used with a learning rate of 2e52e-5 and provides weight decay regularization, which prevents the weights from growing too large to help mitigate overfitting [40]. Overfitting occurs when a model fits too closely to the training dataset, leading to poor performance on the testing data. A scheduler for learning rate decay is also used to reduce the learning rate over the training period. The scheduler gradually increases the learning rate from zero to a specified learning rate during the warmup period, then linearly decreases the learning rate over the remaining epochs of training [41]. The experiments were conducted on NVIDIA GeForce RTX 2080 GPUs. This hardware enabled us to use significant computational power, facilitating quicker processing and learning.
VI-C Embedding Extraction
@@
VI-D Model Configuration
The selected hyperparameters of the model include the number of unique bytes (i.e. 258), hidden size, number of hidden layers, number of attention heads, the intermediate size, maximum length of the input payload, and the number of labels for this application, and are presented in Table II.
The unique bytes parameter represents the total number of unique elements found in the input data. Here, the total is 258, accounting for each unique hexadecimal value and additional pad, unknown, and masking integers, following standard self-supervised practices. The hidden size parameter is essential in determining the dimension of the hidden state. The number of hidden layers represents the number of stacked transformer blocks present in the model. Complementing these blocks is the number of attention heads, which control the model’s ability to attend to different segments of the input. This is pivotal for the model to comprehend and learn from the input data effectively. The intermediate size determines the dimension of the intermediate later in the network within each transformer layer. This value helps ensure that the network can process and transform the input data at each layer effectively. The maximum position embeddings denote the maximum length of the input sequence. For this study, we do not restrict the maximum packet length, which is set to 1500. It is important to note that transformer architectures are inherently capable of processing sequences that extend beyond a single packet’s MTU, meaning that if a payload exceeds 1460 bytes and is split across multiple packets, the transformer can process it as a continuous stream of bytes across packet boundaries. However, increasing the maximum input length significantly raises computational resource requirements, such as memory usage and processing time. These configurations remained consistent through our study of extracting the packet byte embeddings.
TABLE II: Model Configurations
@@ VI-E Downstream Task
TABLE III: Downstream M.L. classifier architecture
Layer Output dim. Components
FC: $d_{\text{model}}\to 512$ 512 ReLU, BatchNorm, Dropout(0.5)
FC: $512\to 128$ 128 ReLU, BatchNorm
FC: $128\to C$ $C$ softmax
In this architecture, shwon in Table III, the input embedding vector of dimension $d_{\text{model}}$ is first passed through a fully connected linear layer that projects it into a 512-dimensional space. This is followed by batch normalization, a ReLU activation to introduce non-linearity, and a dropout layer with a dropout rate of 0.5 to prevent overfitting. The network then reduces the 512-dimensional output to 128 dimensions through a second linear layer, again applying batch normalization and a ReLU activation to maintain stability and non-linearity in the learning process. Finally, a linear layer maps the 128-dimensional vector to $C$ output neurons, where $C$ corresponds to the number of classes. A softmax activation function is applied to the final output to generate a probability distribution over the classes.
@@
VII Few-Shot Learning Architecture
@@
In each episode, we randomly select $K$ classes from the dataset. For each class $k$, two distinct sets are sampled: a support set $S_{k}$ comprising $N_{S}$ examples, and a query set $Q_{k}$ comprising $N_{Q}$ examples. The algorithm computes the class prototypes $\mathbf{c}_{k}$ by taking the mean of the transformed feature vectors of the support set embeddings, $S_{k}$. Once the class prototypes are established, the algorithm evaluates each query sample $(x,y)$ in $Q_{k}$, calculating the Euclidean distance from $x$ to each class prototype $\mathbf{c}_{k}$, and then assigns the class label of the nearest prototype to $x$ . The update to the model parameters is based on the loss $J$, which is formulated as the negative log-likelihood of the softmax probabilities over the distances between the query samples and the class prototypes. Specifically, for each query $(x,y)$ in $Q_{k}$, the loss $J$ is updated according to the equation:
$\scriptsize J\leftarrow J+\frac{1}{KN_{Q}}\left[d(f_{\theta}(x),c_{k})+\log\sum_{k^{\prime}}\exp(-d(f_{\theta}(x),c_{k^{\prime}}))\right].$ (10)
@@
TABLE IV: Results of Binary and Multiclass Classification of Malware Types on the Test Dataset.
Note: Bold font indicates the highest performance, while underlined denotes the second-best result in each column.
@@
VIII Results
In this section, we delve into the metrics to report performance, the state-of-the-art (SOA) benchmarks used for comparison, and the results of our proposed method. We pre-train the encoder once per experimental setting and then run each downstream classifier three times with different random seeds, reporting the mean test accuracy (and episode-averaged mean for FSL).
VIII-A Metrics
@@
VIII-B Benchmarks
TABLE V: Average Inference Time Per Sample (ms)
@@ VIII-B Benchmarks
TABLE VI: Results of Few-Shot Classification of Malware Types on the Query Set
@@
VIII-C Binary and Multi-Class Classification
The data for the downstream classification task involving the SSL embeddings are divided into training, testing, and validation sets, with proportions of 70%, 20%, and 10%, respectively. The performance of our proposed method was evaluated over 25 epochs, demonstrating improvements in accuracy, precision, recall, and F1-Score when using packet bytes as input, shown in Table IV.
@@ VIII-C Binary and Multi-Class Classification
The inference times per sample, shown in Table V, further emphasize the efficiency of the proposed method. The M.L. fully connected networks has a significantly lower inference time compared to other benchmark models. The M.L network achieves an inference time of 0.03 ms and 0.04 ms for UNSW-NB15 and CIC-IOT23, respectively. In contrast, models such as LSTM and CNNs exhibit considerably higher inference times, with the LSTM model requiring over 6 ms per sample on UNSW-NB15. The faster inference time of the M.L. network make it ideal for real-time applications.
@@
VIII-D Few-Shot Classification
VIII-D1 Experimental Setup
The primary motivation for these experiments is to assess the model’s ability to generalize to novel classes with limited supervision, a critical requirement for real-world scenarios where labeled data for all classes may not always be available. FSL allows us to test how well the SSL embeddings can adapt to previously unseen classes with very few labeled examples. To simulate this scenario, we train the SSL model exclusively on two of the three malware classes during pre-training, ensuring no prior exposure to the unseen class.
@@ VIII-D1 Experimental Setup
VIII-D2 Results
The results of our few-shot learning experiments for classifying both known and newly introduced malware types are presented in Table VI. The Trained Classes column indicates which classes were pre-trained with self-supervision, with the Unseen Class column serving as the novel, untrained class for that specific experiment. In the 1-shot setting on the UNSW-NB15 dataset, the model’s highest performance occurs when trained on Exploits and Generic malware types, with an overall query accuracy of 83.04% and 82.82% accuracy specifically for the unseen class. In the 5-shot setting, the best overall query accuracy of 89.48% is achieved when the model is trained on Fuzzers and Generic malware types, with Exploits as the unseen class. Meanwhile, the highest accuracy for an unseen class is observed for the Generic class, achieving 89.95%. For the CIC-IoT23 dataset, the model demonstrates a similar trend but with slightly lower accuracies, reflecting the increased diversity of the IoT traffic data. In the 1-shot setting, the best performance is achieved when training on Backdoor and Brute Force malware types, with Vulnerability as the unseen class, resulting in an overall query accuracy of 62.35% and 61.73% accuracy for the unseen class. Similarly, in the 5-shot setting, the model achieves its highest query accuracy of 75.01% when trained on Backdoor and Brute Force malware types, with Vulnerability as the unseen class.
Figure 6: Few-Shot ROC for Unseen Class on UNSW-NB15 and CIC-IoT23, 1-/5-shot.
@@ VIII-D2 Results
Furthermore, we opt to compare the average few-shot results with the benchmarks to analyze how the accuracy compares, shown in Fig. 5. Despite using significantly fewer training samples, the FSL method performs comparably to SOA classifiers, surpassing methods like 2D-CNN, LSTM, and DEEPShield classifiers. Additionally, in our FSL framework, we train on only two of three classes, unlike the previously mentioned multiclass classification problem, which required all three classes to undergo self-supervised training. This in-turn saves training costs and allows for the ability to detect previously unseen malware classes with high accuracy, an issue that is overlooked in typical classification problems. Overall, our approach effectively leverages knowledge from known malware types to classify novel malware types that were unseen during training, even with a limited number of labeled samples. This capability makes it a promising solution for real-world scenarios where labeled data is scarce and new malware types are frequent.
Lastly, as shown in Fig. 6, we present the Receiver Operating Characteristic (ROC) curves corresponding to the unseen-class few-shot experiments. Each curve demonstrates model performance in both the 1-shot and 5-shot settings, reported separately for the UNSW-NB15 and CIC-IoT23 datasets. By plotting the true positive rate against the false positive rate, these curves highlight the model’s ability to distinguish unseen classes. For UNSW-NB15, all unseen classes achieve strong separability, with AUC values above 0.90 in both 1-shot and 5-shot settings. This indicates that even with minimal labeled support data, the learned embeddings provide highly discriminative features, and performance further improves with additional support examples. For CIC-IoT23, unseen detection is more challenging, as reflected in lower AUC values (0.73–0.78) in the 1-shot setting. However, the curves show improvement with 5-shots, with AUC values improving above 0.86 across all unseen categories. This demonstrates that the model generalizes better when provided with slightly more support samples, particularly in IoT traffic where class distributions and payload patterns are more diverse.
TABLE VII: Masking Percentage Ablation with M.L. Classifier (Binary)
@@
IX Ablation Study
IX-A Masking Percentage
To investigate how varying proportion of masked bytes influences the quality of the learned embeddings, we conduct an ablation study examining multiple masking percentages ranging from 5% to 30%. Although 15% masking is generally used for SSL tasks, our objective is to learn the optimal ratio that preserves sufficient information for robust training of the model, but also prevents the model from becoming too reliant on unmasked data. Table VII displays the results for different masking percentages on distinguishing between malicious and benign packets. On the UNSW-NB15 dataset, we can observe the highest accuracies occur between 5% and 15%, with accuracy slightly dropping off when masking more than 15% of packet bytes. On the more challenging CIC-IoT23 dataset, we observe that higher masking rates improve performance, with 25% byte masking resulting in over a 2% increase in classification over standard 15% masking.